OpenStudy (anonymous):

A skirt costs $20 more than the blouse. The skirt costs $65. What is the price of the blouse. Write an equation. Then solve the problem.

OpenStudy (anonymous):

oh wrong place this is math

OpenStudy (anonymous):

woops i realise this is the wrong question

OpenStudy (anonymous):

Find x if y=1/2 4x/y=2(x+1)